The solar eclipse is indeed a momentous, or at least visually entertaining and curious happening in astrology.

Astronomer Martin Meinshausen claimed that the Dresden Codex, a Mayan manuscript, contained recordings of lunar and solar eclipses.

The Dresden Codex was for the Mayans a way to predict eclipses. 3 Sample Paragraph

From what we have seen concerning this section of the Dresden Codex, it seems likely that Martin Meinshausen proposed that this data was related to the timing of series of solar or lunar eclipses. Our current understanding of the motions of the sun, earth, and moon shows that after the first eclipse in a series, subsequent eclipses can follow at periods of 6 lunar months, approximately 177 or 178 days. At the end of a series of eclipses, a final interval of 5 lunar months, or approximately 148 days, may occur between the last two. The occurrences of exactly these numbers 177 and 148 in the Dresden Codex is the strongest evidence for this interpretation. The lists of three consecutive day names in the 260-day Sacred Round of the Mayan calendar occurring above these numbers may indicate dates on which actual eclipses occurred in the past, or give patterns for predicting future eclipses.